Rodgers and Winter had capital balances of $60,000 and $90,000, respectively, at the beginning of the current fiscal year. The articles of partnership provide for salary allowances of $25,000 and $30,000, respectively; an allowance of interest at 12% on the capital balances at the beginning of the year; and the remaining net income divided equally. Net income for the current year was $110,000.

a. Present the Division of net income section of the income statement for the current year.

Net income $110,000
Rodgers Winter Total
Division of net income:
Salary allowance $ $ $
Interest allowance
Total
Net income $ $ $

b. Assuming that the net income had been $65,000 instead of $110,000, present the Division of net income section of the income statement for the current year.

Net income $65,000
Rodgers Winter Total
Division of net income:
Salary allowance $ $ $
Interest allowance
     Total
Net income $ $ $
Answer #1

Answer:

Ans. A
Net income      1,10,000
Rodgers Winter Total
Division of net income
A Salary allowance         25,000       30,000         55,000
B Interest allowance(60000*12%)(90000*12%)           7,200       10,800         18,000
C Remaining income(37000/2         18,500       18,500         37,000
D=A+B+C Net income         50,700       59,300      1,10,000
*Remaining income = Net income - (salary allowance + interest allowance)
110000 - (55000 + 18000)
110000 - 73000
37000
Ans. B
Net income         65,000
Rodgers Winter Total
Division of net income
A Salary allowance         25,000       30,000         55,000
B Interest allowance           7,200       10,800         18,000
C=A+B Total         32,200       40,800         73,000
D Excess of allowance over net income(8000/2)          -4,000        -4,000          -8,000
E=C-D Net income         28,200       36,800         65,000
CALCULAITON FOR B:
Excess of allowance over net income = Net income - Total allowances
=   65000 - 73000
=-8000